OpenStudy (anonymous):

3x + 2y = 4 2x - y = 5 This system of equations

OpenStudy (anonymous):

From the first equation \[x = \frac{ 4 - 2y }{ 3 }\] Replace x in the second equation \[2*\frac{ 4-2y }{ 3 } - y = 5\] Then you can find y = -1, then replace y with -1 in a previous equation containing x like 3x + 2y = 4. This would become 3x - 2 = 4, therefore x = 2 So x = 2 and y = -1